OpenStudy (anonymous):

Please help - will give you medal (: The following function defines a recursive sequence: f(0) = -5 f(1) = 20 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? (4 points)

OpenStudy (anonymous):

Really will appreciate it (:

OpenStudy (anonymous):

@cwrw238

OpenStudy (cwrw238):

have you got a list of sequences?

OpenStudy (anonymous):

These are the answer choices: -5, -20, -65, -200, … -5, 20, -92, 372, … -5, -24, -92, -372, … -5, 20, -65, 200, …

OpenStudy (anonymous):

idk why there are question marks just ignore them @cwrw238 thanks :)

OpenStudy (cwrw238):

well plug in the value of 2 for n in the formula for f(n) and see what you get - this will give you the thrd term in the sequence ( after 20)

OpenStudy (cwrw238):

then n = 3 for 4th term

OpenStudy (cwrw238):

what terms do the (n-1) and (n-2) refer to when n = 2?

OpenStudy (anonymous):

well if n=2 then the 'n' in those equations will be 2

OpenStudy (cwrw238):

right so (2-1) = which term?

OpenStudy (anonymous):

n? sorry I am a little confused, these aren't my strong suits

OpenStudy (cwrw238):

2-1 = 1 of course and this refers to f(1) which = 20 so which term does (n-2) refer to?

OpenStudy (cwrw238):

n-2 = 2-2 which refers to f(0) which = -5 so these 2 numbers replace (n-1) and (n-2) in the formula

OpenStudy (anonymous):

oh okay I see... @cwrw238 so how would you do that to get the next numbers for the sequence?